OpenStudy (anonymous):

identify if it is a quadratic equation or not 2x^2-root3x+9=0

OpenStudy (anonymous):

\[2x^{2}-\sqrt{3x}+9=0\]

OpenStudy (triciaal):

A quadratic equation is one in which the highest power of the independent variable is 2 genaral format ax^2 + bx + c = 0 where a, b not = 0

Directrix (directrix):

2x² - √(3x) + 9 = y is not the equation of a parabola. If it were a quadratic equation and if 2x² - √(3x) + 9 = 0 were solvable by the Quadratic Formula, then the equation would be that of a Quadratic. as opposed to: 2x² - √(3) x + 9 = y which has a parabola as its graph and is solvable by the Quadratic Formula.